# On the third homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$

### v1

@@ Abstract.
In this article we study the third homology groups of the special linear group $\mathit{{\rm SL}}_{2}(A)$ and the projective special linear group $\mathit{{\rm PSL}}_{2}(A)$, through the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$. As an application we prove a refined Bloch-Wigner exact sequence over certain local domains.
For any commutative ring $A$ and any non-negative integer $n$, the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$ induces the natural homomorphism of homology groups
@@
with 2-power torsion kernel and cokernel. In many problems in algebra and geometry, it is important to understand these kernels and cokernels. In this article we study this map for $n\leq 3$ over certain class of rings. As our first main theorem we prove that if $A$ is a domain, then we have the exact sequence
@@
where $\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is a quotient of $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ over a certain 2-torsion subgroup and $\mu_{2}(A)^{\sim}=1$ if ${\rm char}(A)=2$ and $\mu_{2}(A)^{\sim}\simeq\mathbb{Z}/4$ if ${\rm char}(A)\neq 2$ (see Theorem 2.1 for the precise statement). Moreover, we show that the image of the map $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincides with the image of $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$, where $w={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}\in\mathit{{\rm SL}}_{2}(A)$. Observe that the five term exact sequence from the right of the above exact sequence is the well-known five term exact sequence coming from the central extension
As a particular case, if $A$ is a local domain such that its residue field has more than two elements, then we have the exact sequence
Furthermore, we will study the above 9-term exact sequence over the ring $\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free integer (see Proposition 3.1). As the second main result of this article, we will prove a refined Bloch-Wigner exact sequence over certain local rings. More precisely, let $A$ be a local domain such that either $-1\in({A^{\times}})^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$. Moreover, let $A/\mathfrak{m}_{A}$ be either infinite or it has $p^{d}$ elements where $(p-1)d>6$. Then we prove the refined Bloch-Wigner exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow\mathcal{RB}(A)\rightarrow 0,$
where $\mathcal{RB}(A)$ is the refined Bloch group of $A$ defined by Hutchinson [16] (see Theorem 4.4 for the precise statement). In the above two exact sequences the map
@@
If ${\rm char}(A)=2$, clearly this product map is trivial. We will see that when $A$ is a finite field, a quadratically closed field or a real closed field, then the above product map is trivial. If $A$ is a local domain such that its residue field has more than two elements, then the natural product map
@@
We do not know the answer of the above question even over a general infinite field. At the end of Sections 2 and 3, we give examples of non-local domains that this product map is not trivial.
[…]
@@ 1. The low dimensional homology of central extensions
This gives us a natural embedding
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow\frac{H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})}{\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\ \oplus H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})}.$
@@
(see [29, Proposition 4.4(i), Chap. V]). Thus we get the natural map
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}(\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})).$
@@
The following theorem is proved by Eckmann and Hilton in [12], which generalises Ganea’s 6-term exact sequence [29, Theorem 2.2, Chap. V].
Theorem 1.1.
For any central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$, there is a natural map $\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})$ and a natural exact sequence
@@
Here $\Gamma$ is Whitehead’s quadratic functor from the category of abelian groups to itself. For more on this functor we refer the reader to [4, §1.2] and [19, §2]. Let us explain the map $\tau$. The inclusion $i:A\hookrightarrow G$ induces the natural map $i_{\ast}:\mathcal{A}\rightarrow\mathcal{G}^{\rm ab}=H_{1}(\mathcal{G},\mathbb{Z})$. By the universal property of $\Gamma$, the quadratic map
$\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}),\ \ \ a\mapsto a\otimes i_{\ast}(a),$
can be extended to a group homomorphism
$\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}).$
Since the composite $\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})\overset{\rho_{\ast}}{\longrightarrow}H_{2}(\mathcal{G},\mathbb{Z})$ is trivial, where the second map is given by $a\otimes[g]\mapsto{\boldsymbol{c}}(a,g)$, the composite
[…]
is trivial too. This induces the natural map ${\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z})$. In this article, for calculations purposes, we will use the (left) bar resolution $B_{\bullet}(\mathcal{G})\rightarrow\mathbb{Z}$ of $\mathbb{Z}$ over $\mathbb{Z}[\mathcal{G}]$ to study the integral homology groups $H_{n}(\mathcal{G},\mathbb{Z})$ (see [6, Chap.I, §5]):
If any two of the elements $g_{1},\dots,g_{n}\in G$ commute, then we define
where $S_{n}$ is the symmetric group of degree $n$. For some of properties of this element see [18, Lemma 2.3]. Let $\mathcal{G}$ be a group. Let $K(\mathcal{G},n)$ be an Eilenberg-MacLane space of type $(\mathcal{G},n)$. This is a CW-complex such that
$\pi_{k}(K(\mathcal{G},n),\ast)\simeq\begin{cases}0&\text{if $k\neq n$}\\ \mathcal{G}&\text{if $k=n$}.\end{cases}$
This space is unique up to homotopy. Note that if $n\geq 2$, then $\mathcal{G}$ should be abelian. It is well-known [14, Theorem 21.1] that
$\Gamma(\mathcal{G})\simeq H_{4}(K(\mathcal{G},2),\mathbb{Z}).$
Let $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ be a central extension. The standard classifying space theory gives a (homotopy theoretic) fibration [9, Chap. 6] of Eilenberg-MacLane spaces $K(\mathcal{A},1)\rightarrow K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)$ (see [26, 5.1.28]). From this we obtain the fibration
(see [17, Lemma 3.4.2]). From the Serre spectral sequence of this fibration we obtain the sequence
[…]
This is the sequence
[…]
Let $\mathcal{A}$ be a finite cyclic group. If $2\mid|\mathcal{A}|$, let $\mathcal{A}^{\sim}$ denote the unique non-trivial extension of $\mathcal{A}$ by $\mathbb{Z}/2$. If $2\nmid|\mathcal{A}|$, we define $\mathcal{A}^{\sim}:=\mathcal{A}$. Let $\mathcal{A}$ be an ind-cyclic group, i.e. $\mathcal{A}$ is direct limit of its finite cyclic subgroups. Then by passing to the limit we can define $\mathcal{A}^{\sim}$. Note that if $\mathcal{A}$ has no $2$-torsion, then $\mathcal{A}^{\sim}=\mathcal{A}$. There is always a canonical injective homomorphism $\mathcal{A}\hookrightarrow\mathcal{A}^{\sim}$ whose composition with the projection $\mathcal{A}^{\sim}\rightarrow\mathcal{A}$ coincides with multiplication by $2$.
@@
This follows from the facts that $\Gamma$ commutes with direct limits and $\Gamma(\mathbb{Z}/n)\simeq(\mathbb{Z}/n)^{\sim}$. ∎
Let $\mathcal{A}$ be an abelian group. Let $\sigma_{1}:{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ be obtained by interchanging $\mathcal{A}$. This map is induced by the involution $\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}$, $a\otimes b\mapsto-b\otimes a$ [20, §2]. Let $\Sigma_{2}^{\prime}=\{1,\sigma^{\prime}\}$ be the symmetric group of order 2. Consider the following action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$:
@@
For any abelian group $\mathcal{A}$ we have the exact sequence
@@
If $\mathcal{A}$ is an ind-cyclic group, then the action of Σ2superscriptsubscriptnormal-$\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ is trivial.
Proof.
For the first part see [30, Lemma 5.5] or [5, Section §6]. The second part follows from the first. ∎

### v3

# On the connections between the low dimensional homology groups of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$

@@ Abstract.
In this article we study the low dimensional homology groups of the special linear group $\mathit{{\rm SL}}_{2}(A)$ and the projective special linear group $\mathit{{\rm PSL}}_{2}(A)$, $A$ a domain, through the natural surjective map $\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)$. In particular, we study the connection of the first, the second and the third homology groups of these groups over euclidean domains $\mathbb{Z}[\frac{1}{m}]$, $m$ a square free integer, and local domains.
Let $A$ be a commutative ring with unit and $\mu_{2}(A)$ be the group of $2$-roots of unity in $A$. By studying the Lyndon/Hochschild-Serre spectral sequence associated to the central extension
$1\rightarrow\mu_{2}(A)\rightarrow\mathit{{\rm SL}}_{2}(A)\rightarrow\mathit{{\rm PSL}}_{2}(A)\rightarrow 1,$
and using the homology of cyclic groups [6, pages 58-59] and the universal coefficient theorem [6, Exercise 3, §1, Chap III] it follows that, for any non-negative integer $n$, the natural homomorphism of homology groups
@@
has 2-power torsion kernel and cokernel. In this article we study this map for $n\leq 3$.
As our first main theorem we prove that if $A$ is a domain, then we have the 9-term exact sequence
@@
where $\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ is a quotient of $H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ over a certain 2-torsion subgroup and
$\mu_{2}(A)^{\sim}\simeq\begin{cases}0&\text{ if ${\rm char}(A)=2$}\\ \mathbb{Z}/4&\text{ if ${\rm char}(A)\neq 2$}\end{cases}$
(see Theorem 3.1 for the precise statement). Moreover, if ${\rm char}(A)\neq 2$, we show that the image of the map $\mu_{2}(A)^{\sim}\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ coincides with the image of the map $H_{3}(\langle w\rangle,\mathbb{Z})\rightarrow\widetilde{H}_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$, where $w={\left(\!\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\right)}\in\mathit{{\rm SL}}_{2}(A)$. Observe that the five term exact sequence from the right of the above exact sequence is the well-known five term exact sequence coming from the above central extension [6, Corollary 6.4, Chap. VII].
As first particular case, we study the above 9-term exact sequence over the euclidean domain $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free integer (Proposition 4.1). In particular we show that if $m$ is even, then we have the exact sequence
and if $m$ is odd, then the sequence
is exact.
As second particular case we study the above 9-term exact sequence over local domains (Proposition 5.1). In particular, if the residue field of a local domain $A$ is different than $\mathbb{F}_{2}$, then we obtain the exact sequence
$0\rightarrow\mathbb{Z}/4\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow 0$
We will see that this is not necessarily true if the residue field of the local domain is $\mathbb{F}_{2}$. For example we show that for $\mathbb{Z}_{(2)}=\{a/b\in\mathbb{Q}:a,b\in\mathbb{Z},2\nmid b\}$, we have the exact sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow{\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}{\mu_{2}(\mathbb{Z}_{(2)})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})}}\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}_{(2)}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0.\end{array}$
In all the above exact sequences the map
@@
We will show that for $A_{2}=\mathbb{Z}[\frac{1}{2}]$, $\rho_{\ast}$ is trivial and for $A_{3}=\mathbb{Z}[\frac{1}{3}]$ the map $\rho_{\ast}$ is non-trivial (see Example 4.4).
If $A$ is a local domain such that its residue field has more than two elements, then the natural product map
@@
We do not know the answer of the above question even over a general infinite field. But we show that $\rho_{\ast}$ is trivial for finite fields, quadratically closed fields and real closed fields (Proposition 5.4).
Here we outline the organization of the present paper. In Section 1, we recall some results from the literature over the low dimensional homology groups of central extensions. In Section 2 we introduce the scissors congruence group, the Bloch group and formulate the Bloch-Wigner exact sequence over fields. In Section 3 we prove our 9-term exact sequence (Theorem 3.1). In Section 4, we study the above $9$-term exact sequence over the euclidean domain $\mathbb{Z}[\frac{1}{m}]$. Finally in Section 5 we investigate the above $9$-term exact sequence over local rings.
Notations. In this paper all rings are commutative, except possibly group rings, and have the unit element $1$. For a commutative ring $A$, $\mu(A)$ denotes the group of roots of unity in $A$, i.e.
$\mu(A):=\{a\in A:\text{there is $n\in\mathbb{N}$ such that $a^{n}=1$}\},$
and $\mu_{2}(A):=\{a\in A:a^{2}=1\}$. If $\mathcal{B}\rightarrow\mathcal{A}$ is a homomorphism of abelian groups, by $\mathcal{A}/\mathcal{B}$ we mean ${\rm coker}(\mathcal{B}\rightarrow\mathcal{A})$.
@@ 1. The low dimensional homology of central extensions
If
$\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z}):=\frac{H_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})}{\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})\ \oplus H_{2}(\mathcal{A},\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z})},$
then from the above exact sequence we obtain the exact sequence
$0\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\oplus H_{3}(\mathcal{G},\mathbb{Z})\rightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow 0$
This exact sequence splits naturally, thus we have a natural splitting
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\hookrightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z}).$
@@
(see [26, Proposition 4.4(i), Chap. V]). Thus we get the natural composite map
${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},H_{1}(\mathcal{G},\mathbb{Z}))\rightarrow\widehat{H}_{3}(\mathcal{A}\times\mathcal{G},\mathbb{Z})\rightarrow H_{3}(\mathcal{G},\mathbb{Z})/\rho_{\ast}(\mathcal{A}\otimes_{\mathbb{Z}}H_{2}(\mathcal{G},\mathbb{Z})).$
@@
In this article, for calculations purposes, we will use the (left) bar resolution $B_{\bullet}(\mathcal{G})\rightarrow\mathbb{Z}$ of $\mathbb{Z}$ over $\mathbb{Z}[\mathcal{G}]$ to study the integral homology groups $H_{n}(\mathcal{G},\mathbb{Z})$ (see [6, Chap.I, §5]):
If any pair of the elements $g_{1},\dots,g_{n}\in G$ commute, then we define
where $S_{n}$ is the symmetric group of degree $n$. If $\mathcal{A}:=\langle g_{1},\dots,g_{n}\rangle\subseteq\mathcal{G}$, then the Pontryagin product (see [6, Chap. V, §5]) induces a natural map
$\begin{array}[]{c}\bigwedge_{\mathbb{Z}}^{n}\mathcal{A}\rightarrow H_{n}(\mathcal{A},\mathbb{Z})\rightarrow H_{n}(\mathcal{G},\mathbb{Z}).\end{array}$
It is straightforward to show that the above composite, maps $g_{1}\wedge\cdots\wedge g_{n}$ to ${\boldsymbol{c}}(g_{1},\dots,g_{n})$.
A function $\psi:\mathcal{A}\rightarrow\mathcal{B}$ of (additive) abelian groups is called a quadratic map if
(1) for any $a\in\mathcal{A}$, $\psi(a)=\psi(-a)$,
(2) the function $\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{B}$ with $(a,b)\mapsto\psi(a+b)-\psi(a)-\psi(b)$ is bilinear.
For any abelian group $\mathcal{A}$, there is a universal quadratic map
$\gamma:\mathcal{A}\rightarrow\Gamma(\mathcal{A})$
such that for any quadratic map $\psi:\mathcal{A}\rightarrow\mathcal{B}$, there is a unique group homomorphism $\Psi:\Gamma(\mathcal{A})\rightarrow\mathcal{B}$ such that $\Psi\circ\gamma=\psi$. It is easy to see that $\Gamma$ is a functor from the category of abelian groups to itself. This functor is called Whitehead’s quadratic functor. In fact if $\mathcal{E}$ is the free abelian group generated by the symbols $w(a)$, $a\in\mathcal{A}$, then we can define $\Gamma(\mathcal{A}):=\mathcal{E}/\mathcal{K}$, where $\mathcal{K}$ is the subgroup generated by the relations
(a) $w(a)-w(-a)$=0,
(b) $w(a+b+c)-w(a+b)-w(a+c)-w(b+c)+w(a)+w(b)+w(c)=0$.
The universal quadratic map $\gamma:\mathcal{A}\rightarrow\Gamma(\mathcal{A})$ is given by $a\mapsto\gamma(a):=w(a)+\mathcal{K}$. For more on Whitehead’s quadratic functor we refer the reader to [4, §1.2].
Now let $\mathcal{A}$ be a central subgroup of $\mathcal{G}$. The inclusion $i:\mathcal{A}\rightarrow\mathcal{G}$ induces the natural map $i_{\ast}:\mathcal{A}\rightarrow H_{1}(\mathcal{G},\mathbb{Z})$, $a\mapsto{\boldsymbol{c}}(a)$. By the universal property of $\Gamma$, the quadratic map
$\psi:\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}),\ \ \ a\mapsto a\otimes{\boldsymbol{c}}(a),$
can be extended to a group homomorphism
(1.1) $\tau:\Gamma(\mathcal{A})\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}H_{1}(\mathcal{G},\mathbb{Z}).$
[…]
is trivial too. This induces the natural map
${\rm coker}(\tau)\rightarrow H_{2}(\mathcal{G},\mathbb{Z}).$
Let $\mathcal{G}$ be a group. Let $K(\mathcal{G},n)$ be an Eilenberg-MacLane space of type $(\mathcal{G},n)$. This is a CW-complex such that
$\pi_{k}(K(\mathcal{G},n),\ast)\simeq\begin{cases}1&\text{if $k\neq n$}\\ \mathcal{G}&\text{if $k=n$}.\end{cases}$
This space is unique up to homotopy. Note that for $n\geq 2$, $\mathcal{G}$ should be abelian. It is well-known [14, Theorem 21.1] that
$\Gamma(\mathcal{A})\simeq H_{4}(K(\mathcal{A},2),\mathbb{Z}).$
Let $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$ be a central extension. The standard classifying space theory gives a (homotopy theoretic) fibration [9, Chap. 6] of Eilenberg-MacLane spaces
$K(\mathcal{A},1)\rightarrow K(\mathcal{G},1)\rightarrow K(\mathcal{Q},1)$
(see [23, 5.1.28]). From this we obtain the fibration
(see [16, Lemma 3.4.2]). From the Serre spectral sequence of this fibration we obtain the sequence
[…]
$H_{4}(\mathcal{Q},\mathbb{Z})\rightarrow\ker(\tau).$
The following theorem is proved by Eckmann and Hilton in [12], which generalises Ganea’s 6-term exact sequence [26, Theorem 2.2, Chap. V].
Theorem 1.1 (Eckmann-Hilton).
For any central extension $1\rightarrow\mathcal{A}\rightarrow\mathcal{G}\rightarrow\mathcal{Q}\rightarrow 1$, we have the natural $10$-term exact sequence
@@
Let $\mathcal{A}$ be a finite cyclic group. If $2\mid|\mathcal{A}|$, let $\mathcal{A}^{\sim}$ denote the unique non-trivial extension of $\mathcal{A}$ by $\mathbb{Z}/2$. If $2\nmid|\mathcal{A}|$, we define $\mathcal{A}^{\sim}:=\mathcal{A}$. Thus if $n=|\mathcal{A}|$, then
$\mathcal{A}^{\sim}\simeq\begin{cases}\mathbb{Z}/2n&\text{if $2\mid n$}\\ \mathbb{Z}/n&\text{if $2\nmid n$.}\end{cases}$
Let $\mathcal{A}$ be an ind-cyclic group, i.e. $\mathcal{A}$ is direct limit of its finite cyclic subgroups. Then by passing to the limit we can define $\mathcal{A}^{\sim}$. Note that if $\mathcal{A}$ has no $2$-torsion, then $\mathcal{A}^{\sim}=\mathcal{A}$. There is always a canonical injective homomorphism $\mathcal{A}\hookrightarrow\mathcal{A}^{\sim}$ whose composition with the projection $\mathcal{A}^{\sim}\rightarrow\mathcal{A}$ coincides with multiplication by $2$.
@@
This follows from the facts that the functor $\Gamma$ commutes with direct limit and
$\Gamma(\mathbb{Z}/n)\simeq\mathbb{Z}/{\rm gcd}(n^{2},2n)\simeq(\mathbb{Z}/n)^{\sim},$
where ${\rm gcd}(n^{2},2n)$ is the greatest common divisor of $n^{2}$ and $2n$ (see [4, Corollary 1.2.9]). ∎
Let $\mathcal{A}$ be an abelian group. Let $\sigma_{1}:{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$ be obtained by interchanging the copies of $\mathcal{A}$. This map is induced by the involution $\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathbb{Z}}\mathcal{A}$, $a\otimes b\mapsto-b\otimes a$ [18, §2]. Let $\Sigma_{2}^{\prime}=\{1,\sigma^{\prime}\}$ be the symmetric group of order 2 and consider the following action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})$:
@@
Let $\mathcal{A}$ be an abelian group and $T_{\mathcal{A}}$ be its torsion subgroup. Then we have the exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})^{\Sigma_{2}^{\prime}}\rightarrow 0.\end{array}$
If $T_{\mathcal{A}}$ is an ind-cyclic group, then $\Sigma_{2}^{\prime}$ acts trivially on ${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$ and the exact sequence
$\begin{array}[]{c}0\rightarrow\bigwedge_{\mathbb{Z}}^{3}\mathcal{A}\rightarrow H_{3}(\mathcal{A},\mathbb{Z})\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})\rightarrow 0\end{array}$
splits naturally.
Proof.
By [27, Lemma 5.5] or [5, Section §6] we have the exact sequence
@@
Since ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathcal{A},\mathcal{A})\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$, we obtain the first exact sequence. Now let $T_{\mathcal{A}}$ be an ind-cyclic group. Since ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathbb{Z}/n,\mathbb{Z}/n)\simeq\mathbb{Z}/n$, the action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(\mathbb{Z}/n,\mathbb{Z}/n)$ is trivial. Now by passing to the limit, we see that the action of $\Sigma_{2}^{\prime}$ on ${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})$ is trivial. For the last part note that since $T_{\mathcal{A}}$ is ind-cyclic, $\bigwedge_{\mathbb{Z}}^{3}T_{\mathcal{A}}=0$. Now applying the first part to the inclusion $T_{\mathcal{A}}\hookrightarrow\mathcal{A}$, we obtain the following commutative diagram with exact rows
[…]
Now from this diagram we obtain a natural splitting map
${{\rm Tor}_{1}^{\mathbb{Z}}}(T_{\mathcal{A}},T_{\mathcal{A}})\simeq H_{3}(T_{\mathcal{A}},\mathbb{Z})\rightarrow H_{3}(\mathcal{A},\mathbb{Z}).$